OpenStudy (anonymous):

Point E Is located At (3,8) And Point T Is Located At (12,3). What Is The Length Of ET? Round your answer to the nearest tenth

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Hi, You want to use the distance formula. What exactly is the distance formula? It is used to find the units between two points. \[d = \sqrt{(x _{2} - x _{1}})^2 + (y _{2} - y _{1})^2\] In here, we have the points: [(3, 18), (12, 3)] Where: x2 = 12 x1 = 3 y2 = 3 y1 = 18. Plug these into the Distance Formula and solve. OpenStudy Ambassador; Compassionate.

OpenStudy (anonymous):

Very nicely done. Note that the square root extends all the way over to include the y values.